OpenStudy (anonymous):

What is the sum of the geometric sequence 1, 3, 9, … if there are 12 terms? 292,524 265,720 139,968 104,976

OpenStudy (anonymous):

ooh easy

OpenStudy (anonymous):

ar^0+ar^1+ar^2... +ar^11 the rule 1 2 3 ... 12 term 1 3 9 ? result a=1 r=3 (1-3^12)/(1-3)=265720 ^ here before i had 12 lol. It was just a typo

OpenStudy (anonymous):

found a and r this way. since ar^0= a =1 a is 1 so we look at 2nd term ar^1=3 it just reduces to r=3 now we use the geometric series formula http://www.purplemath.com/modules/series5.htm (1-3^12)/(1-12)=48312.727272 <- with the error